OpenStudy (anonymous):

find the slope if it exists the containg the pair of points (-5.7,18.7 ) and (-5.3,3.1) the slope is?

OpenStudy (anonymous):

formula is m=(y2-y1)/(x2-x1) (3.1-18.7)/(-5.3 - -5.7) = -15.6/0.4 =-39